# Turbulence and far-from-equilibrium equation of state of Bogoliubov waves in Bose-Einstein Condensates

### v2

@@ Abstract
Now we will use the short-wave solution (12) for interpreting the experimental results of Dogra et al. Dogra et al. (2023) on the far-from-equilibrium EoS. Dogra et al. analyzed the particle spectra (occupation number spectra) $\tilde{n}_{k}$ measured in the direct cascade of BEC turbulence by fitting a scaling-law of $k^{-3+\delta}$, with $\delta$ being a small adjustable parameter. Note that at small scales, our prediction (12) corresponds to a $k^{-3}$-wave-action spectrum (equivalent to the particle spectrum, in this case, up to a prefactor difference), which coincides, up to the log-correction, with the 4-wave prediction found in Zhu et al. (2023a). Such a coincidence makes both spectra good candidates for explaining the experimentally found EoS.
@@
$\tilde{n}_{k}=n_{\rm 4w}k^{-3}\log^{-\frac{1}{3}}\left(\frac{k}{k_{\rm f}}\right)\,,n_{\rm 4w}=C_{\rm 4w}\left(\frac{m^{2}\epsilon}{\hbar^{3}a_{s}^{2}}\right)^{1/3}\,,$ (13)
@@
where $k_{\rm f}$ is the forcing wave number and $C_{\rm 4w}={C_{d}}{(4\pi)^{-\frac{2}{3}}}$, with $C_{\rm d}\approx 5.26\times 10^{-2}$ an universal constant. $m$ is the atom mass, $\hbar$ is the reduced Planck constant, and $a_{s}$ is the $s$-wave length. The energy flux $\epsilon$ is defined as the energy per unit of volume dissipated in the system per unit time.
@@
In Reference Dogra et al. (2023), the EoS is defined as the dependence of the non-dimensional spectrum amplitude, $n_{4w}/N$ vs the non-dimensional energy flux $\tilde{\epsilon}=m^{2}\hbar^{-3}a_{s}^{-2}N^{-3}\epsilon$ where $N$ is the mean particle density such that $N=\int 4\pi k^{2}\tilde{n}_{k}\,\mathrm{d}k$ with the definition of $\tilde{n}_{k}$ as in this paper. Theoretical prediction (13) implies $n_{4w}/N=C_{\rm 4w}\tilde{\epsilon}^{1/3}$, which was confirmed in Dogra et al. (2023) for data points with low values of $\tilde{\epsilon}$. However, at larger $\tilde{\epsilon}$, Dogra et al. (2023) reported a steeper EoS, $n_{4w}/N\sim\tilde{\epsilon}^{\alpha}$ with $\alpha>1/3$. The authors of Dogra et al. (2023) noted that $\alpha\approx 2/3$ provide a good fit implying a flux-dependence like the one of the classical Kolmogorov hydrodynamic turbulence. However, it is unclear how a state with the classical Kolmogorov turbulence could appear in this experiment, considering that it is unlikely that the cascade process is generated by vortices with hydrodynamic properties. Indeed, such regimes are usually attributed to states with polarised vortex tangles at scales greater than the inter-vortex distance Polanco et al. , which do not seem to be relevant to the considered experiment. Also, as noted in Dogra et al. (2023), the Kolmogorov spectrum would imply a spectral exponent that is different from the observed exponent $\approx-3$. Another point is that the EoS observed in Dogra et al. (2023) shows a dependence on the mean particle density $N$ that can not be explained by (13).
@@
Here, we suggest a plausible interpretation of the steeper EoS and its dependence on $N$. We propose that a stronger ground-state condensate component could be present for greater $\tilde{\epsilon}$ making the system to switch to the three-wave turbulence of short Bogoliubov waves, implying $\alpha\approx 1/2>1/3$.
Recasting our prediction (9) in the same manner as in (13), we obtain (see Appendix D for details)
$\tilde{n}_{k}=n_{\rm 3w}k^{-3}\,,\frac{n_{\rm 3w}}{N}=C_{\rm 3w}\left(\frac{m^{2}\epsilon}{\hbar^{3}N^{3}a_{s}^{2}}\right)^{1/2}\,,C_{\rm 3w}=\frac{C_{2}}{2^{17/4}\pi^{2}}\,.$ (14)
@@
The dimensional $P_{0}$ in (9) is related to $\epsilon$ as $\epsilon=\rho_{0}P_{0}$, since $P_{0}$ complies with the standard definition of the energy flux as the energy dissipation rate per unit mass whereas $\epsilon$ used in Dogra et al. (2023) is the energy dissipation rate per unit volume. The EoS corresponding to (14) is $n_{\rm 3w}/N=C_{\rm 3w}\tilde{\epsilon}^{1/2}$.
Figure 2: EoS generated from the experimental data of Dogra et al. (2023). Green and magenta points correspond to the spectra that are best fitted by the three-wave and the four-wave predictions respectively.
The revised EoS based on the experimental data points of Dogra et al. (2023) is shown in Fig. 2. In Dogra et al. (2023), the EoS is presented as a dependence $n_{0}/N$ vs $\tilde{\epsilon}$, with $n_{0}$ found from fitting the experimental spectra with $\tilde{n}_{k}=n_{0}k^{-3.2}$. Our procedure is different: for each experimentally obtained spectrum, we fit our two theoretical predictions $\tilde{n}_{k}=n_{\rm 4w}^{\rm exp}k^{-3}\,\log^{-\frac{1}{3}}\left(\frac{k}{k_{\rm f}}\right)$, and $\tilde{n}_{k}=n_{\rm 3w}^{\rm exp}k^{-3}$. We then select the spectrum amplitude to be either $n_{\rm 4w}^{\rm exp}$ or $n_{\rm 3w}^{\rm exp}$ depending on which one provides a better fit (see Appendix D). Figure 2 also shows the theoretical four-wave and three-wave EoS predictions in which the constants $C_{\rm 4w}$ and $C_{\rm 3w}$ are replaced by the mean values of the respective constants obtained by fits of the experimental spectra, $\langle C_{\rm 4w}^{\rm exp}\rangle$ and $\langle C_{\rm 3w}^{\rm exp}\rangle$. We see that the theoretical predictions are consistent with the experimental data, with the four-wave and three-wave EoS in agreement with the low and high $\tilde{\epsilon}$ parts respectively. The mean experimental constants are nearly twice the theoretically predicted values, $\langle C_{\rm 3w}^{\rm exp}\rangle\approx 2.2C_{\rm 3w}$ and $\langle C_{\rm 4w}^{\rm exp}\rangle\approx 2.4C_{\rm 4w}$. However, the spread of the EoS data points appears to be approximately the same, $\Delta C_{\rm 4w}\approx 2.3C_{\rm 3w}$ and $\Delta C_{\rm 4w}\approx 1.7C_{\rm 4w}$, which are of a similar size to the spread of the experimental points. We conclude that our theoretical predictions agree with the experimental data within the measurement error bars.
@@
In summary, in this Letter, we obtained the Kolmogorov-type stationary spectrum of forced-dissipated turbulence of short Bogoliubov waves, Eq.(12), and we revised the long-wave Kolmogorov-type spectrum including finding the exact prefactor constant $C_{1}$, see (9). We have validated these solutions, along with the combined spectrum including both long and short waves by numerical simulations of the GPE and the isotropic Bogoliubov WKE. We used the newly found short-wave spectrum for interpreting the (previously unexplained) experimental Equation of State of Ref.Dogra et al. (2023). To test our theory in future experiments, the realizability of short Bogoliubov waves must be examined, namely (i) check if a strong ground state is present for strong forcing cases, (ii) check that for these cases the measured spectrum corresponds to waves shorter than the healing length and with amplitudes weaker than the ground state. Further, it would be interesting to study stronger waves whose statistics is not described by the WTT.

### v4

@@ Abstract
Almost all BEC experiments relevant to wave turbulence have been interpreted within the framework of the 4-wave interaction theory Navon et al. (2016); Dogra et al. (2023); Moreno-Armijos et al. (2025), which assumes that the wave perturbations $\delta\Psi$ greatly exceed the condensate amplitude $\Psi_{0}$, leading to cubic nonlinearity. Instead, in this work we employ the new short-wave solution (3) governed by 3-wave interactions to interpret the experimental results of Dogra et al. Dogra et al. (2023) on the far-from-equilibrium EoS. Dogra et al. analyzed the particle spectra (occupation number spectra) $n_{k}$ (which has a difference of prefactor of $8\pi^{3}$ from the one we defined as (18) in Methods) measured in the direct cascade of BEC turbulence by fitting a scaling-law of $k^{-3+\delta}$, with $\delta$ being an adjustable parameter. Note that at small scales, our prediction (3) corresponds to a $k^{-3}$ wave-action spectrum (equivalent to the particle spectrum, in this case, up to a prefactor difference; see (A16) in Methods), which coincides, up to the log-correction, with the 4-wave prediction found in Zhu et al. (2023). This coincidence makes both spectra good candidates for explaining the experimentally found EoS, assuming that they correspond to the short-wave range such that $k\xi>1$.
@@
$n_{k}=n_{\rm 4W}\,k^{-3}\log^{-\frac{1}{3}}\left(\frac{k}{k_{\rm f}}\right)\,,n_{\rm 4W}=C_{\rm 4W}\left(\frac{m^{2}\epsilon}{\hbar^{3}a_{s}^{2}}\right)^{\frac{1}{3}}\,$, (13)
@@
where $k_{\rm f}$ is the forcing wave number and $C_{\rm 4W}={C_{d}}{(4\pi)^{-\frac{2}{3}}}$, with $C_{\rm d}\approx 5.26\times 10^{-2}$ being a universal constant. Here, $m$ is the atom’s mass, $\hbar$ is the reduced Planck constant, and $a_{s}$ is the $s$-wave length. The energy flux $\epsilon$ is defined as the energy dissipated in the system per unit time per unit volume.
In Dogra et al. (2023), the authors performed a series of experiments with various parameters, interaction strength $a_{s}$, forcing intensity, and mean particle density $N$ such that $N=\int 4\pi k^{2}n_{k}\,\mathrm{d}k$ with the definition of $n_{k}$ as in this paper. They obtained the particle spectra amplitude $n^{\rm exp}$ by fitting the experimental spectra with $n_{k}=n^{\rm exp}k^{-3.2}$, and discovered that employing the dimensional variables $n^{\rm exp}$ and $\epsilon$ as two state variables results in a non-universal EoS for different $a_{s}$ and $N$. Furthermore, they empirically found that the dependence of the non-dimensional spectrum amplitude, $n_{\rm exp}/N$ versus the non-dimensional energy flux $\tilde{\epsilon}=m^{2}\hbar^{-3}a_{s}^{-2}N^{-3}\epsilon$ collapses the experimental data into a universal EoS. Theoretical prediction (13) implies
$n_{4W}/N=C_{\rm 4W}\tilde{\epsilon}^{1/3},$ (14)
@@
which was confirmed in Dogra et al. (2023) for data points with low values of $\tilde{\epsilon}$. However, at larger $\tilde{\epsilon}$, Dogra et al. (2023) reported a steeper EoS, $n_{4W}/N\sim\tilde{\epsilon}^{\alpha}$ with $\alpha>1/3$. The authors of Dogra et al. (2023) noted that $\alpha\approx 2/3$ provide a good fit implying a flux-dependence like the one of the classical Kolmogorov hydrodynamic turbulence. However, it is unclear how a state with the classical Kolmogorov turbulence could appear in this experiment, considering that it is unlikely that the cascade process is generated by vortices with hydrodynamic properties. Indeed, such regimes are usually attributed to states with polarised vortex tangles at scales greater than the inter-vortex distance Polanco et al. (2021), which do not seem to be relevant to the considered experiment. Also, as noted in Dogra et al. (2023), the Kolmogorov spectrum would imply a spectral exponent 5/3-5/3 that is very different from the observed exponent $\approx-3$.
@@
Here, we suggest a plausible interpretation of the steeper EoS. We propose that a stronger ground-state condensate component could be present for greater $\tilde{\epsilon}$ making the system switch to the 3-wave turbulence of short Bogoliubov waves, implying $\alpha\approx 1/2>1/3$.
Recasting our prediction (3) in the same manner as in (13), we obtain (Methods)
$n_{k}=n_{\rm 3W}k^{-3}\,,\frac{n_{\rm 3W}}{N}=C_{\rm 3W}\left(\frac{m^{2}\epsilon}{\hbar^{3}N^{3}a_{s}^{2}}\right)^{\frac{1}{2}}\,,C_{\rm 3W}=\frac{C_{2}}{2^{17/4}\pi^{2}}\,.$ (15)
@@
The dimensional $P_{0}$ in (2) is related to $\epsilon$ as $\epsilon=\rho_{0}P_{0}$, since $P_{0}$ complies with the standard definition of the energy flux as the energy dissipation rate per unit mass whereas $\epsilon$ used in Dogra et al. (2023) is the energy dissipation rate per unit volume. The EoS corresponding to (15) is
$n_{\rm 3W}/N=C_{\rm 3W}\tilde{\epsilon}^{1/2}.$ (16)
This EoS looks similar to the 4-wave EoS (14): the both laws are of the form $n/N=C\,\tilde{\epsilon}^{\alpha}$. Hovewer, one should not forget the important difference between the 4-wave and the 3-wave case: the non-normalised 4-wave EoS given by (13) does not contain any $N$-dependence whereas the 3-wave law given in (15) does involve $N$.
Figure 2: Equation of State (EoS) regenerated from the experimental data of Dogra et al. (2023). (a) Particle spectrum amplitude $n^{\rm exp}$ .vs. energy flux $\epsilon$, where $n^{\rm exp}$ is taken either as $n^{\rm exp}_{4W}$ in green fitting by the 4-wave prediction (13) or $n^{\rm exp}_{3W}$ in magenta fitting by the 3-wave prediction (15). The fitting procedure is indicated by the inset of (a): Each experimental spectrum $n^{\textbf{exp}}_{k}$ is compensated by $k^{3}\log^{1/3}(\frac{k}{k_{\rm f}})$ and $k^{3}$, respectively. $n^{\rm exp}_{4W}$ and $n^{\rm exp}_{4W}$ are calculated by averaging the corresponding compensated spectrum on a plateau range. We take $n^{\rm exp}$, $n^{\rm exp}_{4W}$ or $n^{\rm exp}_{3W}$ depending on which is closer to its theoretical prediction, $n_{4W}$ or $n_{3W}$. (b) Non-dimensional particle spectrum amplitude $n^{\rm exp}/N$ .vs. non-dimensional flux $\tilde{\epsilon}$. Theoretical predictions $C_{\rm 4W}\tilde{\epsilon}^{1/3}$ and $C_{\rm 3W}\tilde{\epsilon}^{1/2}$ and the ones replacing $C_{\rm 4W}$ and $C_{\rm 3W}$ by the mean values of the respective constants obtained by fits of the experimental spectra, $\langle C_{\rm 4W}^{\rm exp}\rangle$ and $\langle C_{\rm 3W}^{\rm exp}\rangle$, are superimposed.
Re-analysis of the experimental data
The revised EoS based on the experimental data points of Dogra et al. (2023) is shown in Fig. 2. Our procedure is different from the one of Dogra et al. (2023): for each experimentally obtained spectrum $n^{\text{exp}}_{k}$, we try our two theoretical predictions, namely $n_{k}=n_{\rm 4W}^{\rm exp}k^{-3}\,\log^{-\frac{1}{3}}\left(\frac{k}{k_{\rm f}}\right)$, and $n_{k}=n_{\rm 3W}^{\rm exp}k^{-3}$, investigate the compensated spectra $n^{\text{exp}}_{k}k^{3}\log^{1/3}(\frac{k}{k_{\rm f}})$ and $n^{\text{exp}}_{k}k^{3}$, respectively, and compute $n^{\rm exp}_{\rm 4W}$ and $n^{\rm exp}_{\rm 3W}$ by averaging the corresponding compensated spectrum in their plateau ranges, as shown in the inset of Fig. 2(a). We take the value of the spectrum amplitude, $n^{\rm exp}$, either $n^{\rm exp}_{\rm 4W}$ or $n^{\rm exp}_{\rm 3W}$, depending on which is closer to its corresponding prediction, $n_{\rm 4W}$ or $n_{\rm 3W}$. In the other words, we find the best fit for the compensated spectra in terms of the $L^{1}$-norm. We plot the fitting spectrum amplitude $n^{\rm exp}$ vs. the flux $\epsilon$ in Fig. 2(a), magenta for the points taken with $n^{\rm exp}_{\rm 4W}$ and green for those taken with $n^{\rm exp}_{\rm 3W}$. The non-dimensional EoS is presented in Fig. 2(b), in which the points split into two distinct clouds and collapse onto a universal EoS. Note that in Dogra et al. (2023), the authors reduced the spread of their EoS by introducing two extra phenomenological fitting parameters in addition to $n^{\rm exp}/N$ and $\tilde{\epsilon}$. Our fitting procedure is based only on comparison with our two theoretical predictions without any introducing any adjustable parameters.
Figure 2 (b) also shows the theoretical 4-wave and 3-wave EoS predictions in which the constants $C_{\rm 4W}$ and $C_{\rm 3W}$ are replaced by the mean values of the respective constants obtained by fits of the experimental spectra, $\langle C_{\rm 4W}^{\rm exp}\rangle$ and $\langle C_{\rm 3W}^{\rm exp}\rangle$. We see that the theoretical predictions are consistent with the experimental data, with the 4-wave and 3-wave EoS in agreement with the low and high $\tilde{\epsilon}$ parts respectively. The mean experimental constants are nearly twice the theoretically predicted values, $\langle C_{\rm 3W}^{\rm exp}\rangle\approx 2.2C_{\rm 3W}$ and $\langle C_{\rm 4W}^{\rm exp}\rangle\approx 2.4C_{\rm 4W}$. The deviation of the experimentally measured constants from their theoretical predictions is similar of the same order as the spread ($L^{\infty}$-norm, $\Delta C^{\rm exp}=C^{\rm exp}_{\rm max}-C^{\rm exp}_{\rm min}$) of the EoS data points from the mean, $\Delta C^{\rm exp}_{\rm 4W}\approx 2.3C_{\rm 3W}$ and $\Delta C^{\rm exp}_{\rm 4W}\approx 1.7C_{\rm 4W}$. Since there are no adjustable parameters in our theoretical results and fitting process, and theoretical predictions are asymptotical for $k\xi\gg 1$ (which in experiments are not satisfied in a strong sense, as $k\xi\approx 4$) we conclude that our theoretical predictions reasonably agree with the experimental data. Our numerical simulations with shaken traps, which are more realistic in terms of corresponding to the experimental setup, show that the interaction between the cloud and trap can significantly increase the measured constants of the stationary KZ-type spectrum, which is also observed in the previous study of Zhu et al. (2023). This agrees with the observation that $\langle C_{\rm 4W}^{\rm exp}\rangle>C_{\rm 4W}$ and $\langle C_{\rm 3W}^{\rm exp}\rangle>C_{\rm 3W}$. We expect that in larger traps, the experimental measurements would become closer to the universal limit and their measured constants would tend towards the predicted values. Nevertheless, further studies of the inhomogeneity induced by traps and respective systematic deviations in numerical (and experimental) flux measurements are desirable to better comprehend the finite size effects.
Discussion
@@
In summary, in this paper we have obtained the Kolmogorov-type stationary spectrum of forced-dissipated turbulence of short Bogoliubov waves, (3), and we revised the long-wave Kolmogorov-type spectrum including finding the exact prefactor constant $C_{1}$, see (2). We have validated these solutions, along with the combined spectrum including both long and short waves by numerical simulations of the GPE and the isotropic Bogoliubov WKE. Our work relies on the turbulence universality hypothesis, stating that the inertial (non-dissipative) range spectrum is insensitive to specific forcing and dissipation mechanisms. Consequently, we adopted stochastic forcing and hyperviscous dissipation to maximize the inertial range, as is customary in turbulence simulations. Our prediction were also tested in more realistic settings, by using an oscillating external trap to force the system. We used the newly found short-wave spectrum for interpreting the (previously unexplained) experimental Equation of State of Ref.Dogra et al. (2023). To test our theory in future experiments, the realizability of short Bogoliubov waves must be examined, namely (i) check if a strong ground state is present for strong forcing cases, (ii) check that for these cases the measured spectrum corresponds to waves shorter than the healing length and with amplitudes weaker than the ground state. It is natural to expect that the short waves of a real system are not in an ideal regime where $|\Psi_{0}|\gg|\delta\Psi|$, so that the 3-wave interaction prevails, or a regime where $|\Psi_{0}|\ll|\delta\Psi|$, so that the 4-wave interaction dominats. This may contribute to further variance in constant values between our predictions and the experimental results. A future model is anticipated to describe short waves when the condensate amplitude and fluctuation amplitude are comparable, such that $|\Psi_{0}|\sim|\delta\Psi|$. This regime corresponds to stronger waves whose statistics is not described by the WTT.
Although the non-dispersive acoustic ZS spectrum is not the main topic of this work, it is an interesting side result that we would like to comment on. We find a way to regularize the WKE of non-dispersive acoustic waves for the 3D isotropic case, but get a different prefactor. Consequently, the ZS spectrum comes with a bigger constant (Methods). In recent work Kochurin and Kuznetsov (2024), the authors numerically obtained the ZS spectra in 3D acoustic wave systems, and different values of constant were reported for the dispersive and non-dispersive cases. The constant observed in Kochurin and Kuznetsov (2024) for the non-dispersive acoustic system is smaller than the one for the dispersive system, unlike what we get here. Their paper offers an explanation based on a numerical observation of an anisotropy (i.e. presumably the isotropic spectrum is unstable with respect to anisotropic perturbations). Note, however, that the model acoustic-wave equation studied in Kochurin and Kuznetsov (2024) was different from the compressible Euler equation or GPE. Thus, the question about the acoustic wave spectrum arising in the non-dispersive Euler equation merits further study.
Methods
Wave-kinetic equations and steady Kolmogorov-type spectrum
The starting point to derive the 3-wave interaction amplitude $V^{1}_{23}$ in (8) is to rewrite the Hamiltonian of the second order $H_{2}$ into the canonical form of $H_{2}=\Sigma_{{\bf k}}\omega_{\bf k}a_{\bf k}a^{*}_{\bf k}$, then the wave-action spectrum is defined as
$n^{\text{B}}_{\bf k}=\left(\frac{L}{2\pi}\right)^{3}\langle|a_{\bf k}|^{2}\rangle\,$, (17)
where the problem is considered in an $L$-periodic box. As a consequence, one rewrites the Hamiltonian of third order into $H_{3}=\sum_{1,2,3}V^{1}_{2,3}\left(a_{1}a^{*}_{2}a^{*}_{3}+\text{c.c}\right)\delta^{1}_{2,3}\,$ and gets the interaction amplitude. In the Supporting Information (SI), we provide two methods to derive $V^{1}_{23}$. Just to mention one of the two methods here, we first assume $\Psi(\bm{x},t)=\Psi_{0}(1+\psi(\bm{x},t))$, such that $|\Psi_{0}|^{2}=\rho_{0}$, $\langle\psi(\bm{x},t)\rangle=0$ and $|\psi|\ll 1$, and defines the Fourier series as $\psi(x)=\Sigma_{\bf k}\psi_{\bf k}e^{i{\bf k}\cdot{\bf x}}$. We obtain $a_{\bf k}$ from $\psi_{\bf k}$ by finding a linear Bogoliubov transformation, and this allows us to find the relationship between the particle spectrum (ignoring the zero mode), defined as
$n_{\bf k}=\left(\frac{L}{2\pi}\right)^{3}\langle|\Psi_{\bf k}|^{2}\rangle/m=\left(\frac{L}{2\pi}\right)^{3}\langle|\psi_{\bf k}|^{2}\rangle\rho_{0}/m\,$, (18)
and the wave-action spectrum $n^{\text{B}}_{\bf k}$. One should note that we define the particle spectrum so that $N=\int 4\pi k^{2}n_{k}\mathrm{d}k$, where $N$ is the mean particle density. A different convention was used in Dogra et al. (2023): the particle spectrum was defined in the way that $N=\frac{1}{(2\pi)^{3}}\int 4\pi k^{2}n_{k}\mathrm{d}k$. This difference in notation has to be taken into account when dealing with the experimental data.
In the acoustic limit, $\psi_{\bf k}\to\frac{1}{2\xi\sqrt{c_{\text{s}}k}}(a_{\bf k}-a^{*}_{-{\bf k}})$. Assuming that $a_{\bf k}$ and $a_{-{\bf k}}$ have the same amplitude and independent random phases (consistent with the so-called random amplitude and phases assumption in WWT plus the isotropy assumption), one gets
$n_{\bf k}=\frac{\rho_{0}}{2m\xi^{2}c_{\text{s}}k}n^{\text{B}}_{\bf k}\,.$ (19)
In the short-wave limit, we have $\psi_{\bf k}\to\frac{1}{\sqrt{\sqrt{2}c_{\text{s}}\xi}}a_{\bf k}$, which results in
$n_{\bf k}=\frac{\rho_{0}}{\sqrt{2}mc_{\text{s}}\xi}n^{\text{B}}_{\bf k}\,.$ (20)
To obtain isotropic WKE, we calculate the angular average of the Dirac-$\delta$ function of wavevectors in (7), $\Delta=\langle\delta({\bf k}-{\bf k}_{1}-{\bf k}_{2})\rangle$ in SI, yielding $\Delta=\frac{2\pi}{kk_{1}k_{2}}\,$. Remarkably, we discover that for the degenerate resonant manifold of non-dispersive sound waves, $\Delta_{\rm nondisp}=\frac{\pi}{kk_{1}k_{2}}$, which regularizes the 3D non-dispersive isotropic WKE of acoustic turbulence. Nonetheless, to make sense of the WKE, one should take the limit of the large box first, followed by the angle average, and finally the large time limit (small nonlinearity). The resulting isotropic 3D acoustic WKE has the same form as (10), except that the prefactor is now $2\pi^{2}V_{0}^{2}/c_{\text{s}}$. In SI, we present a revised, rigorous derivation of the WKE (9), and the parametrization of the resonant manifold defined by (5).
We seek for stationary power-law solutions of wave-kinetic equations using the standard process that the WTT offers Nazarenko (2011). We provide all the technical information in SI, including the calculation and locality analysis, while the main text outlines the essential steps. Note that for the non-dispersive acoustic case, we have $P(k)=8\pi^{3}V_{0}^{2}A^{2}\frac{I(x)}{2x-9}k^{9-2x}$, in which the non-dimensional collision integral $I(x)$ is the same as the one for dispersive acoustic waves, and obtain
$C_{\rm 1,nondisp}=\frac{\sqrt{2}}{\sqrt{3\pi(4\ln{2}+\pi-1)}}\,.$ (21)
run model $N_{r}$ $k_{\text{min}}$ $k_{\text{max}}$ $k_{f}$ $s$
1 Bogoliubov WKE 3276832768 $1\times 10^{-2}/\xi$ $200/\xi$ 0 4
2 acoustic WKE 8196 $\times 10^{-4} 11$ 0.01 2
run model $f_{0}$ $\Delta k_{f}$ $\beta$ $k_{d}$
1 Bogoliubov WKE $1.032\times 10^{14}$ $1.5\times 10^{-2}$ 1 80
2 acoustic WKE $2.419\times 10^{16}$ 0.028 1 0.6
run model $\xi$ $N_{p}$ $\alpha$ $k_{R}$ $f_{\rm w}^{2}$ $\gamma$
3 GPE $1.25\Delta x$ 960 3 260 $0^{-9} 1-1$
4 GPE $15\Delta x$ 576 4 300 $0^{-4} 1-1$
5 GPE $1.25\Delta x$ 384 3 128
6 GPE $10\Delta x$ 384 4 200
Table 1: Parameters for GPE and WKE simulations.
Numerical simulation methods
We perform GPE simulations using the standard massively-parallel pseudo-spectral code FROST Krstulovic (2020) with a fourth-order Exponential Runge-Kutta temporal scheme Zhu et al. (2022). We use grids of $N_{p}^{3}$ collocation points so that the grid length is $\Delta x=L/N_{p}$. The forced-dissipated GPE (4) is used for the GPE runs 3 and 4, with stochastic forcing term supported on a narrow band ($1\leq k\leq 3$) and obeying the Ornstein–Ulenbeck process $\mathrm{d}\mathcal{F}_{\bf k}(t)=-\gamma\,\widehat{\Psi}_{\bf{k}}\mathrm{d}t+f_{\rm w}\mathrm{d}{W}_{\bf k}$ where $W_{\bf k}$ is the Wiener process. The parameters $\gamma$ and $f_{0}$ control the correlation time and the amplitude of the forcing, respectively. In addition, the condensate amplitude (at $k=0$) is kept constant during evolution. Dissipation is of the form $D_{\bf k}=(k/k_{\rm R})^{2\alpha}$ acting at the small scales. Finally, the $k$-space energy fluxes $P(k)$ are computed directly using the GPE (4) (see Supplemental Materials of Griffin et al. (2022)). We verify that the simulated WT is in a weakly nonlinear regime by measuring spatio-temporal spectra of $\Psi({\bf x},t)$ and confirming that they are narrowly concentrated around the dispersion curve (1) (see SI).
To simulate BEC with a potential trap, we set a box trap and release the constraint on the $k=0$ mode. The potential term reads $V(\bm{x})=0$ for coordinates satisfying $|x|,|y|,|z|\leq L_{\rm trap}$, and $V_{\rm box}(\bm{x})=U_{D}$ otherwise. $U_{D}=1000\mu$ is the value we select for the GPE runs 5 and 6. $\mu=mc_{s}^{2}$ is the chemical potential. We obtain the ground state using the imaginary-time scheme with this box trap. We then add energy by replacing the forcing term $i\mathcal{F}$ in (4) by an shaking potential term $V_{osc}({\bf x},t)=A\,x\sin(\omega_{\rm res}t)+A\,y\cos(\frac{\sqrt{5}}{2}\omega_{\rm res}t)$. Note that we shake the trap in both $x$ and $y$ directions to isotropize the system, which is particularly necessary to obtain acoustic waves at relatively large scales. The shaking amplitude is set as $A=0.5\mu/L_{\rm trap}$ and the frequency is set as the Bogoliubov frequency of the trap, which is $\omega_{\rm res}=c_{s}k_{\rm trap}\sqrt{1+(k_{\rm trap}\xi)^{2}/2}$. Dissipation is set in the same way as in the periodic case, but the condensate amplitude is not fixed. The energy flux is estimated by $P_{0}=2\int D_{k}E(k)\mathrm{d}k$, and the density, sound speed, and healing length used in Fig 1(a) for runs 5 and 6 are the effective values inside the trap when the system gets stationary.
We simulate isotropic WKE using the code WavKinS Krstulovic and Labarre (2024). Similarly to the GPE, we also include a forcing term of the form $f_{k}=f_{0}k^{s}e^{-(k-k_{\rm f})^{2}/\Delta k_{\rm f}^{2}}$ and dissipation term $(k/k_{d})^{\beta}n_{k}$ in the right-hand side of WKE (9) and (10). The code uses a logarithmic mesh $\{k_{i}=C\lambda^{i},i=1,\ldots,N_{r}\}$, where $\lambda$ is chosen in order to span the integration domain $k_{\rm min},k_{\rm max}]$. Inter-mesh values are obtained using a linear interpolation that ensures positivity. The collision integral is a 1D integral obtained by making explicit use of the parameterization of the resonant manifold provided in SI. Integrals are performed using the trapezoidal rule, after performing a change of variable in the following manner $\int\limits_{k_{i}}^{k_{i+1}}G(p)\mathrm{d}p\approx\frac{\log{\lambda}}{2}\left(G(k_{i+1})k_{i+1}+G(k_{i})k_{i}\right)$. We make an approximation of $\int\limits_{0}^{k_{1}}G(p)\mathrm{d}p\approx k_{1}G(k_{1})$ for the integral on the first bin. Finally, we use a Runge-Kutta-2 time marching method, with an implicit scheme for the dissipative term.
All essential numerical parameters for the WKE and GPE runs are given in Table S1. The choice of parameters in each run were made to maximize the inertial (non-dissipative) range of $k$ in the regime at which this specific run is focused: large-scale acoustic, small-scale Bogoliubov, or the regime containing both large- and small-scale WT.
EoS predicted by 4-wave and 3-wave WT and fitting method
In reference Zhu et al. (2023), the constant energy flux solution in the $4$-wave regime was derived analytically and confirmed numerically. Its $\log$-corrected spectrum reads
$n_{k}=C_{\mathrm{d}}\left(\epsilon\hbar/g^{2}\right)^{1/3}k^{-3}\ln^{-1/3}(k/k_{\rm f})\,.$ (22)
Substituting $g=4\pi\hbar^{2}a_{s}/m$ into the above equation, one gets the equation of states of (13).
We recall the relationship between the particle spectrum and the wave-action spectrum for short Bogoliubov waves (A16), and substitute $n^{\text{B}}_{k}=\frac{E(k)}{4\pi k^{2}\omega_{\rm short}(k)}$, the KZ solution (3), and $P_{0}=\epsilon/\rho_{0}$ into it. This gives
$n_{k}=\frac{C_{2}}{4\pi}\frac{\rho_{0}^{1/2}\xi^{1/2}}{mc_{\text{s}}^{3/2}}\epsilon^{1/2}k^{-3}\,.$ (23)
$n_{k}=\frac{C_{2}}{2^{17/4}\pi^{2}}\frac{m\epsilon^{1/2}}{\hbar^{3/2}N^{1/2}a_{s}}k^{-3}\,$, (24)
and consequently the EoS (15).
In generating Fig. 2, each experimental spectrum is fitted using the 4-wave prediction $n_{k}=n_{\rm 4W}^{\rm exp}k^{-3}\log^{-\frac{1}{3}}\left(\frac{k}{k_{\rm f}}\right)$ and the 3-wave predictions $n_{k}=n_{\rm 3W}^{\rm exp}k^{-3}$, respectively, yielding the experimentally obtained spectrum amplitude $n^{\rm exp}_{\rm 4W}$ and $n^{\rm exp}_{\rm 3W}$. To determine the most suitable candidate between these two models, theoretical predictions are calculated $n_{\rm 4W}$ and $n_{\rm 3W}$ using (13) and (15). The model is then selected based on the computed ratio $n^{\rm exp}_{4W}/n_{\rm 4W}$ and $n^{\rm exp}_{3W}/n_{\rm 3W}$, prioritizing the one whose ratio is closes to unity.